OpenStudy (beaubrun):

N2 + 3H2 → 2HN4 If you produce 55.5 grams of ammonia, how many grams of nitrogen will you need?

OpenStudy (aaronq):

Stoichiometry When performing stoichiometric calculations the values to be used \(\sf \color{red}{must~be~in~moles}\)*. This is because the Molar Masses (\(\sf M_{m}\) of two elements are not the same. This is to say that 1 gram of Hydrogen atoms is not the same number of atoms as 1 gram of Nitrogen atoms. We need to be working with numbers of particles (e.g. atoms, molecules, formula units) and using moles is the way this is done. General Scheme: \(\sf \large 1.\)First write and balance an equation for the process described. \(\sf \large 2.\)Next, use the stoichiometric coefficients to find moles produced. \(\sf \large 3.\) Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\sf \large \color{red}{a}A + \color{blue}{b}B \rightleftharpoons \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\sf \dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\sf\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) \(\sf \large 4.\) Solve

OpenStudy (aaronq):

To interconvert between mass and moles, use the relationship: \(\Large \sf n=\dfrac{m}{M_m}\) where, \(M_m\)=molar mass, \(m\)=mass, and \(n\)= moles.